# Misty: UI Prototyping Through Interactive Conceptual Blending

### v2

@@ Abstract.
journalyear: 2024
@@ 1. Introduction
While it’s unsurprising that other research has demonstrated the utility of conceptual blending in computational creativity support for various creative domains (Câmara Pereira, 2007; Zbikowski, 2018), there are nevertheless limited tools for blending user interfaces, and developers often have to manually implement these blends from scratch. How then, can we bring conceptual blending into our tools for user interface creation?
Towards addressing this gap, we created a web prototype system, Misty [^1] Misty is an acronym for Mixing Interfaces by Semantic Transformation. Plus Y, because why not., that embodies and operationalizes conceptual blending. With Misty, users can upload example UI screenshots or hand-drawn sketches, specify specific regions and aspects of interest, blend them into their work-in-progress UIs, and use a semantic diff to make additional refinements to the blend (Figure 1). We postulated that because generative AI models encode certain characteristics of human ingenuity, these latest multi-modal AI models would have within them some capacity to automatically interpret and apply conceptual blending.
To evaluate Misty, we conducted user studies with 14 frontend developers. The study results revealed three key insights: First, participants found Misty’s varied levels of input granularity valuable, supporting different stages of the design process from early ideation to detailed refinement. Second, while the current AI models showed limitations in generating perfect conceptual blends, participants appreciated our semantic diff and the code editor for tweaking generation results, often using imperfections as springboards for further ideation. Third and finally, conceptual blending through Misty enhanced UI prototyping workflows by supporting both focused iteration on single designs and broad exploration of diverse alternatives.
@@ 2. Example Usage Scenario For Misty
Fuji has recently installed Misty, which streamlines the incorporation of design examples through conceptual blending. With some inspiring UI designs found online, Fuji is ready to use Misty to improve their profile UI code.
@@
Fuji began by uploading the code for their current profile UI to Misty, along with screenshots of a few inspiring profile UI designs. To get a quick overview of possibilities, Fuji used Misty’s global blending feature. They connected the example UI 1 to their source UI 2, triggering a global, screen-level blending operation. Misty generated updated code that took Fuji’s current profile UI content and blended into it the visual styles of the example UI 3.
@@
Misty provided a list of changes alongside the outputs 4, summarizing the code updates. Fuji noticed that Misty helped to turn their UI into dark theme and adopt the list item layout in the example. Fuji decided to keep the updated list item layout, but turn off the dark theme. By switching off the toggle corresponding to dark theme, Fuji received an updated result from Misty 5, which returned the UI back to the previous light theme but kept the other changes.
Usually, Fuji would manually re-create the example using code, which is very time-consuming. Moreover, without seeing their own content reflected in these design variations, Fuji would struggle to get a direct impression of how each design would work in their scenario. But now, with Misty, Fuji can get a quick visual preview of multiple design possibilities, all populated with their own content.
@@ 3.2. Conceptual Blending
Previous studies have demonstrated conceptual blending’s utility in computational creativity support for various creative domains (Câmara Pereira, 2007; Zbikowski, 2018). In music, conceptual blending has been applied to inspire music production by blending unique concepts drawn from music and language, two very distinct domains (Zbikowski, 2018). Conceptual blending has also been used as the basis for studying creativity and building computational implementations of creative acts (Câmara Pereira, 2007). Tools such as CreativeConnect have established blending as an effective operation in creative exploration for graphic design (Choi et al., 2024).
Conceptual blending is intrinsic to the inspiration-seeking processes in UI prototyping (Herring et al., 2009; Lu et al., 2024a). Current practices such as moodboarding also reflects conceptual blending’s value for exploration and ideation (Chen et al., 2019; Ritchie et al., 2011). In this work, we apply conceptual blending to UI development. We built our system Misty on the insight that the latest multi-modal AI models are capable of interpreting user intents, writing UI code (Wu et al., 2024; Si et al., 2024), and have the potential to operationalize conceptual blending because they encode certain characteristics of human ingenuity—for example, in metaphor reasoning (Tong et al., 2024; Lakoff, 2014) and conceptual blending of pop culture images (Wang et al., 2023).
@@ 3.3. AI Support for UI/UX Prototyping
For UI development, large multi-modal AI models have exhibited capabilities in code generation given UI screenshots (Si et al., 2024) or natural language descriptions (Wu et al., 2024; Lu et al., 2023; Li et al., 2023b). While techniques for code generation from natural language have advanced rapidly in recent years (Yin and Neubig, 2017; Le et al., 2020; Li et al., 2023a), generation based on provided UI design screenshots remains limited to simplistic outputs (Si et al., 2024).
@@ 4.1. System Design Goals
Misty embodies three design goals supported by findings from a literature review of three key areas: UI/UX design processes and tools (Kumar et al., 2011; Wu et al., 2023; Warner et al., 2023; Lu et al., 2022; Knearem et al., 2023; Lu et al., 2024a; Kim et al., 2022; Peng et al., 2024), AI for creativity support (Choi et al., 2024; Chung and Adar, 2023; Angert et al., 2023; Chen et al., 2024), and Human-AI Interaction (Horvitz, 1999; Masson et al., 2024; Shneiderman and Maes, 1997; Subramonyam et al., 2023; Amershi et al., 2019). We formulate the design goals as follows:
@@ 2. DG2
    Preserve developer agency in creative interaction. Maintaining designer agency is crucial in AI-assisted design tools (Lu et al., 2022; Knearem et al., 2023; Li et al., 2024). By offering granular control over the blending process, we ensure that the tool remains predictable and aligned with the user’s vision, addressing key concerns raised in previous studies on AI adoption in design workflows (Lu et al., 2022; Knearem et al., 2023). Studies have also shown that users prefer visual interactions over purely textual prompts when working with AI tools (Subramonyam et al., 2023; Masson et al., 2024; Lu et al., 2024a).
@@ 3. DG3
    Promote exploration, reflection, and iteration. Creative design processes are inherently iterative and reflective (Buxton, 2010; Norman, 2013), so conceptual blending should support reflection by allowing the user to visually compare different blended results and to critically evaluate choices before making further changes. This iterative approach aligns with established design practices such as parallel prototyping (Choi et al., 2024; Dow et al., 2010; Peng et al., 2024) that enable users to refine ideas to arrive at more innovative solutions.
@@ 4.2. Interaction Affordances
4.2.1. Conceptual blending control for implicit inte
@@ 4.3. Implementation Details
Misty is implemented in React, React Flow, Tailwind CSS, and React Live, with OpenAI’s GPT-4o as the underlying large multimodal model servicing prompt requests. We employ few-shot prompting (Brown, 2020) with chain-of-thought reasoning (Wei et al., 2022) to improve response quality (see Section A.3 for full details on the prompts used). To detect and repair malformed React syntax, Misty employs heuristic rules implemented using regular expressions. However, in cases where these heuristics are insufficient, we note that users may opt to either re-generate the code or use the code editor to apply manual repairs.
@@ 1. RQ1
    How do developers utilize the different features provided by Misty?
@@ 5.1. Participants
Since Misty uses frontend code to represent UIs, we recruited frontend engineers as our participants. Thus, they were able to fully use all features of Misty, especially the code editor, and understand the limitations of current AI models in generating frontend code (Si et al., 2024). All participants had experience creating and improving UIs by themselves, rather than only implementing designs created by designer colleagues (see more discussion in Section 4).
@@ 5.2. Procedure
We divided our study into the following three steps. When participants performed the tasks, we asked them to think aloud to provide insights into their cognitive process [^3] Section A.1 and A.2 provide a detailed list of questions in our interview and usability questionnaire. (Ericsson and Simon, 1980). We also piloted the study twice to refine the process and validate the tasks were realistic and engaging.
@@ 5.2.3. Step 3 (15 minutes): Usability questionnaire and follow-up interview
After completing the tasks, participants completed a questionnaire on their experience using Misty. We adapted the questionnaire from the standard System Usability Scale (Brooke, 1996) and included five-point Likert scale questions evaluating Misty on its usability, user control and agency, output quality, value for design exploration, and real-world adoption potential. After completing the questionnaire, we conducted a semi-structured interview with participants. The interview questions covered scenarios where participants would or would not use Misty, desired improvements or additional features, concerns about using Misty, and thoughts on how tools like Misty might influence their work.
@@ 6.2.2. Allowing flexible intent specification
Participants appreciated Misty’s diverse input mechanisms for specifying design intent, which supported various stages of their development process. The ability to blend in full screenshots, specific aspects of example UIs, and hand-drawn sketches proved efficient for different phases of UI development. For early ideation, participants valued the option to start exploring with generally blending in full screens or hand-drawn sketches. Particularly, hand-drawn sketches are widely used in roughly laying out design concepts and communicating design ideas. The ability to blend in sketches fits well with developers current practices (P3, P6, P11, P12), especially “in the case (when) I have an idea in mind, for which I cannot find any relevant template online” (P6). P3 noted “a lot of my time is just translating a sketch… this (Misty) can save me time on making the code or the visual look, so I don’t have to fight with CSS and layout”. As designs progressed, participants found value in having more specific input options. P3 noted, for example, that drag-and-drop is very good for updating existing products, instead of starting brand new ones.
@@ 6.2.4. Inspiring through serendipity
However, serendipity can also be a double-edged sword. Although some of the blends initially have an “amazing wow factor” (P10), when carefully scrutinizing the blend there are often small details that were initially overlooked, such as the “separation between items and not apply the button design style to one of the buttons” (P10). The randomness, which “can be good in exploration is also bad when you already know where to go” (P7) and “have a specific idea in mind” (P3). For such scenarios, having a “more deterministic way of working with the tool would be preferred” (P3). This experience made P3 “impressed by some things and yet underwhelmed by others.”
@@ 6.2.5. Refining iteratively and visually
While Misty helps in initial explorations, participants needed more support for iterative visual refinement after the initial exploratory phases (P1, P2, P8–10). Participants often found themselves wanting to make precise adjustments directly over the output, without drag-and-drop of another example. Participants wanted the ability to “add text prompt to the output, like I could when I was adding parts of a design to full styles” (P8). This comes from a demand for pixel-perfect UI implementations in professional contexts, which current models are not always able to produce (Si et al., 2024; Wu et al., 2024). From many UI developers, they are expected to produce pixel perfect implementations, especially by designers (P3). As a result, when seeing issues in the blending outputs’ visual details (e.g. element sizing, alignment, and color values), developers naturally wanted to fix them with direct prompt iterations over the visually rendered UI. This is currently only supported by direct editing in our code editor. However, developers also recognized that they are fine with manually tweaking these details. While the blending outputs were not pixel perfect, developers generally thought they were “good enough output for me to build upon” (P4), then, “the minute interaction of copying over font or changing some spacing, that’s something I can do way quicker by myself” (P14). Yet, developers can still benefit from direct iterations on the visual outputs in Misty, without necessarily involving more examples.
@@ 6.2.6. Blending boundaries between designers and developers
P2, P4, and P13 considered both designers and developers potential user groups that can benefit from using Misty. The system showed potential in blending designer and developer activities in UI prototyping, shifting collaboration and handoff points in the full prototyping workflow. This is noted by eight participants in our study. P1, P4, and P14 mentioned the shifting handoff points between designers and developers, where traditionally they only receive design artifacts that contain no code implementation. If designers use Misty to generate some code, developers can directly take the code and immediately use it in their implementation (P1). Misty can also help developers to incorporate stakeholder feedback faster, enhancing cross-function collaboration (P12) and reducing the friction with designers (P3, P13).
At a higher level, Misty also complements designers and developers with each others’ skills. P2 noted that Misty “does not replace me, but does a good job of covering for my lack of design skills.” Misty helps to raise the bar of design for developers’ UI implementations. Participants imagined designers using Misty may not have to rely on engineers and vice versa (P11). In the future, tools like Misty will continuously lower the boundary of entry for both UI design and development, changing the separated dynamics and goals for these two roles.
@@ 7. Discussion
Artists and creators have expressed concern over content ownership as generative AI models advance (Srinivasan and Uchino, 2021; Flick and Worrall, 2022). Conceptual blending, while prevalent in creative tasks, is essentially borrowing content and styles from existing examples that users do not necessarily own. Conceptual blending tools have to be responsibly deployed with careful ethical considerations around ownership, permissions, and their usages in blending derivatives.
UI/UX design and development pose unique challenges for these ethical considerations in conceptual blending. Since software users do not adapt quickly to new interface and interaction patterns, designers often need to incorporate established patterns that users are already familiar with (Norman, 2013). As a result, usability—instead of novelty or creativity—is often prioritized in UI/UX prototyping. The boundary of appropriation in UI/UX design is often ambiguous, as designers and developers open-source numerous component libraries for others to reuse and re-purpose. One participant during our user study even mentioned that they “use and abuse the web” to directly copy other websites’ HTML structure and CSS attributes as a common part of their workflow.
@@
These are only some challenges that the generative UI community is grappling with that would need to be resolved in this evolving landscape of Generative AI technology.
@@
Our implementation of Misty is based on simple UI screens with limited interactivity. During our user study, participants expressed interest in expanding Misty to conceptual blending of interactive prototypes. Interactive UI prototypes contain many dimensions that are relevant for conceptual blending: user interactions, animations, and screen states. Tools like CoCapture (Chen et al., 2021) have made progress towards lending interactivity, and Keyframer (Tseng et al., 2024) has investigated the use of Generative AI for authoring animations. However, blending interactive continues to pose design challenges. Blending static screenshots and sketches are relatively simple, yet it is not obvious how to specify sections and aspects—for example, of a video—for blending. Design principles from research in programming by demonstration can provide a useful starting point (Cypher and Halbert, 1993).

### v3

@@ Abstract.
UI prototyping, UX design, conceptual blending, artificial intelligence
copyright: rightsretainedjournalyear: 2025conference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713924isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Systems and tools for interaction designccs: Human-centered computing User interface programming
@@ 1. Introduction
While it’s unsurprising that other research has demonstrated the utility of conceptual blending in computational creativity support for various domains (Câmara Pereira, 2007; Zbikowski, 2018), there are nevertheless limited tools for blending user interfaces. Blending visual inspirations is common in UI creation, but designers and developers often have to manually implement these blends from scratch (Herring et al., 2009). How then, can we bring conceptual blending into our tools for user interface creation?
Towards addressing this gap, we created a web prototype system, Misty [^1] Misty is an acronym for Mixing Interfaces by Semantic Transformation. Plus Y, because why not., that embodies and operationalizes conceptual blending. With Misty, users can upload example UI screenshots or hand-drawn sketches, specify specific regions and aspects of interest, blend them into their work-in-progress UIs, and use a semantic diff to make additional refinements to the blend (Figure 1). We postulated that because generative AI models encode certain characteristics of human ingenuity (Lakoff, 2014; Tong et al., 2024), these latest multi-modal AI models would have within them some capacity to automatically interpret and apply conceptual blending. In Misty, we evaluated conceptual blending in developing individual mobile UI screens.
To evaluate Misty, we conducted a user study with 14 frontend developers. The study results revealed three key insights: First, participants found Misty’s varied levels of input granularity valuable, supporting different stages of the design process from early ideation to detailed refinement. Second, while the current AI models showed limitations in generating perfect conceptual blends, participants appreciated our semantic diff and the code editor to tweak generation results, often using imperfections as springboards for further ideation. Third and finally, conceptual blending through Misty enhanced UI prototyping workflows by supporting both focused iteration on single designs and broad exploration of diverse alternatives.
@@ 2. Example Usage Scenario For Misty
Fuji has recently installed Misty, which streamlines the incorporation of design examples through conceptual blending. With some inspiring UI designs found online, Fuji is ready to use Misty to explore improvement directions for the profile UI.
@@
Fuji began by uploading the code for their current profile UI to Misty, along with screenshots of a few inspiring profile UI designs. To get an overview of possibilities, Fuji used Misty’s global blending feature to quickly explore alternatives. They connected the example UI 1 to their source UI 2, triggering a global, screen-level blending operation. Misty generated updated code that took Fuji’s current profile UI content and blended into it the visual styles of the example UI 3.
@@
Misty provided a list of changes alongside the outputs 4, summarizing the code updates. Fuji noticed that Misty helped to turn their UI into dark theme and adopt the list item layout in the example. Fuji decided to keep the updated list item layout, but turn off the dark theme. By switching off the toggle corresponding to dark theme, Fuji received an updated result from Misty 5, which returned the UI back to the previous light theme but kept the other changes. This gave Fuji more control in exploration, by picking the desired categories of change and directly seeing the visual result.
Usually, Fuji would manually re-create the example using code, which is very time-consuming. Moreover, without seeing their own content reflected in these design variations, Fuji would struggle to get a direct impression of how each design would work in their scenario. But now, with Misty, Fuji can get a quick visual preview of multiple design possibilities, all populated with their own content. Misty’s global blending feature supports a more comprehensive exploration of the design space.
@@ 3.2. Conceptual Blending
Previous studies have demonstrated conceptual blending’s utility in computational creativity support for various creative domains (Câmara Pereira, 2007; Zbikowski, 2018). In music, conceptual blending has been applied to inspire music production by blending unique concepts drawn from music and language, two very distinct domains (Zbikowski, 2018). Conceptual blending has also been used as the basis for studying creativity and building computational implementations of creative acts (Câmara Pereira, 2007). Tools such as CreativeConnect have established blending as an effective operation in creative exploration for graphic design (Choi et al., 2024). Researchers have also designed direct manipulation interfaces for AI to support creative analogical reasoning (Srinivasan and Chan, 2024).
Conceptual blending is intrinsic to the inspiration-seeking processes in UI prototyping (Herring et al., 2009; Lu et al., 2024a). Current practices such as moodboarding also reflects conceptual blending’s value for exploration and ideation (Chen et al., 2019; Ritchie et al., 2011). In this work, we apply conceptual blending to UI development. We built our system Misty on the insight that the latest multi-modal AI models are capable of interpreting user intents, writing UI code (Wu et al., 2024; Si et al., 2025), and have the potential to operationalize conceptual blending because they encode certain characteristics of human ingenuity—for example, in metaphor reasoning (Tong et al., 2024; Lakoff, 2014) and conceptual blending of pop culture images (Wang et al., 2023b).
@@ 3.3. AI Support for UI/UX Prototyping
For UI development, large multi-modal AI models have exhibited capabilities in code generation given UI screenshots (Si et al., 2025) or natural language descriptions (Wu et al., 2024; Lu et al., 2023; Li et al., 2023b). While techniques for code generation from natural language have advanced rapidly in recent years (Yin and Neubig, 2017; Le et al., 2020; Li et al., 2023a), generation based on provided UI design screenshots remains limited to simplistic outputs (Si et al., 2025).
@@ 4.1. System Design Goals
Misty embodies three design goals supported by findings from a literature review of three key areas: UI/UX design processes and tools (Kumar et al., 2011; Wu et al., 2023b; Warner et al., 2023; Lu et al., 2022; Knearem et al., 2023; Lu et al., 2024a; Kim et al., 2022; Peng et al., 2024), AI for creativity support (Choi et al., 2024; Chung and Adar, 2023; Angert et al., 2023; Chen et al., 2024), and Human-AI Interaction (Horvitz, 1999; Masson et al., 2024; Shneiderman and Maes, 1997; Subramonyam et al., 2024; Amershi et al., 2019). We formulate the design goals as follows:
@@ 2. DG2
    Preserve developer agency in creative interaction. Maintaining designer agency is crucial in AI-assisted design tools (Lu et al., 2022; Knearem et al., 2023; Li et al., 2024). By offering granular control over the blending process, we ensure that the tool remains predictable and aligned with the user’s vision, addressing key concerns raised in previous studies on AI adoption in design workflows (Lu et al., 2022; Knearem et al., 2023). Studies have also shown that users prefer visual interactions over purely textual prompts when working with AI tools (Subramonyam et al., 2024; Masson et al., 2024; Lu et al., 2024a).
@@ 3. DG3
    Promote exploration, reflection, and iteration. Creative design processes are inherently iterative and reflective (Buxton, 2010; Norman, 2013), so conceptual blending should support reflection by allowing the user to visually compare different blended results and to critically evaluate choices before making further changes. This iterative approach aligns with established design practices, such as parallel prototyping (Choi et al., 2024; Dow et al., 2010; Peng et al., 2024), that help users refine ideas to arrive at more innovative solutions.
@@ 4.2. Interaction Affordances
4.2.1. Conceptual blending control for implicit intent disambiguation
@@ 4.3. Implementation Details
Misty is implemented in React, React Flow, Tailwind CSS, and React Live, with OpenAI’s GPT-4o as the underlying large multimodal model servicing prompt requests. We employ few-shot prompting (Brown et al., 2020) with chain-of-thought reasoning (Wei et al., 2022) to improve response quality (see Section A.3 for full details on the prompts used). To detect and repair malformed React syntax, Misty employs heuristic rules implemented using regular expressions. However, in cases where these heuristics are insufficient, we note that users may opt to either re-generate the code or use the code editor to apply manual repairs.
@@ 1. RQ1
    How do developers utilize the different features in Misty?
@@ 5.1. Participants
Since Misty uses frontend code to represent UIs, we recruited frontend engineers as our participants. Thus, they were able to fully use all features of Misty, especially the code editor, and understand the limitations of current AI models in generating frontend code (Si et al., 2025). All participants had experience creating and improving UIs by themselves, rather than only implementing designs created by designer colleagues (see more discussion in Section 4).
@@ 5.2. Procedure
We divided our study into the following three steps. When participants performed the tasks, we asked them to think aloud to provide insights into their cognitive process [^3] Section A.1 and A.2 provide a detailed list of questions in our interview and usability questionnaire. (Ericsson and Simon, 1980). We also piloted the study twice to refine the process and validate the tasks were realistic and engaging. These two pilot studies were not included in the above 14 participants.
@@ 5.2.3. Step 3 (15 minutes): Usability questionnaire and follow-up interview
After completing the tasks, participants completed a questionnaire on their experience using Misty. We adapted the questionnaire from the standard System Usability Scale (Brooke, 1996) and included 5-point Likert scale questions evaluating Misty on its usability, user control and agency, output quality, value for design exploration, and real-world adoption potential. After completing the questionnaire, we conducted a semi-structured interview with participants. The interview questions covered scenarios where participants would or would not use Misty, desired improvements or additional features, concerns about using Misty, and thoughts on how tools like Misty might influence their work.
@@ 6.1. Quantitative Results
For most participants who rated low on “Misty supports me well in expressing my intents”, they requested features to continuously improve Misty’s output after conceptual blending. For example, after receiving an output, participants wanted to ask the AI model to further iterate on output details. These iterations do not involve reference examples, and are therefore not directly related to our focus on conceptual blending.
When specifically asked about Misty’s features for intent expressions, such as cropping and drag and drop, participants did not raise any issues. They felt these features addressed their needs in blending operations, as we will discuss later in Section 6.2.2.
@@ 6.2.2. Allowing flexible intent specification
Participants appreciated Misty’s diverse input mechanisms for specifying design intent, which supported various stages of their development process. The ability to blend in full screenshots, specific aspects of example UIs, and hand-drawn sketches proved efficient for different phases of UI development. For early ideation, participants valued the option to start exploring with generally blending in full screens or hand-drawn sketches. Particularly, hand-drawn sketches are widely used in roughly laying out design concepts and communicating design ideas. The ability to blend in sketches fits well with developers’ current practices (P3, P6, P11, P12), especially “in the case (when) I have an idea in mind, for which I cannot find any relevant template online” (P6). P3 noted “a lot of my time is just translating a sketch… this (Misty) can save me time on making the code or the visual look, so I don’t have to fight with CSS and layout”. As designs progressed, participants found value in having more specific input options. P3 noted, for example, that drag-and-drop is very good for updating existing products, instead of starting brand new ones.
@@ 6.2.4. Inspiring through serendipity
However, serendipity can also be a double-edged sword. Although some of the blends initially have an “amazing wow factor” (P10), when carefully scrutinizing the blend there are often small details that were initially overlooked, such as the “separation between items and not apply the button design style to one of the buttons” (P10). The randomness, which “can be good in exploration” (P7), is “also bad when you already know where to go” (P7) or “have a specific idea in mind” (P3). For such scenarios, having a “more deterministic way of working with the tool would be preferred” (P3). This experience made P3 “impressed by some things and yet underwhelmed by others.”
@@ 6.2.5. Refining iteratively and visually
While Misty helps in initial explorations, participants needed more support for iterative visual refinement after the initial exploratory phases (P1, P2, P8–10). Participants often found themselves wanting to make precise adjustments directly over the output, without drag-and-drop of another example. Participants wanted the ability to “add text prompt to the output, like I could when I was adding parts of a design to full styles” (P8). This comes from a demand for pixel-perfect UI implementations in professional contexts, which current models are not always able to produce (Si et al., 2025; Wu et al., 2024). From many UI developers, they are expected to produce pixel perfect implementations, especially by designers (P3). As a result, when seeing issues in the blending outputs’ visual details (e.g. element sizing, alignment, and color values), developers naturally wanted to fix them with direct prompt iterations over the visually rendered UI. This is currently only supported by direct editing in our code editor. However, developers also recognized that they are fine with manually tweaking these details. While the blending outputs were not pixel perfect, developers generally thought they were “good enough output for me to build upon” (P4), then, “the minute interaction of copying over font or changing some spacing, that’s something I can do way quicker by myself” (P14). Yet, developers can still benefit from direct iterations on the visual outputs in Misty, without necessarily involving more examples.
@@ 6.2.6. Blending boundaries between designers and developers
P2, P4, and P13 considered both designers and developers as potential user groups that can benefit from using Misty. The system showed potential in blending designer and developer activities in UI prototyping, shifting collaboration and handoff points in the full prototyping workflow. This is noted by eight participants in our study. P1, P4, and P14 mentioned the shifting handoff points between designers and developers, where traditionally they only receive design artifacts that contain no code implementation. If designers use Misty to generate some code, developers can directly take the code and immediately use it in their implementation (P1). Misty can also help developers to incorporate stakeholder feedback faster, enhancing cross-function collaboration (P12) and reducing the friction with designers (P3, P13).
At a higher level, Misty also complements designers and developers with each others’ skills. P2 noted that Misty “does not replace me, but does a good job of covering for my lack of design skills.” Misty helps to raise the bar for design for developers’ UI implementations. Participants imagined designers using Misty may not have to rely on engineers and vice versa (P11). In the future, tools like Misty will continuously lower the boundary of entry for both UI design and development, changing the separated dynamics and goals for these two roles.
@@ 7. Discussion
Artists and creators have expressed concern over content ownership as generative AI models advance (Srinivasan and Uchino, 2021; Flick and Worrall, 2022). Conceptual blending, while prevalent in creative tasks, is essentially borrowing content and styles from existing examples that users do not necessarily own. Conceptual blending tools have to be responsibly deployed with careful ethical considerations around ownership, permissions, and their use in blending derivatives.
UI/UX design and development pose unique challenges for these ethical considerations in conceptual blending. Since software users do not adapt quickly to new interface and interaction patterns, designers often need to incorporate established patterns that users are already familiar with (Norman, 2013). As a result, usability—instead of novelty or creativity—is often prioritized in UI/UX prototyping. For this reason, the boundary of unethical appropriation in UI/UX design is often ambiguous. Moreover, designers and developers open-source component libraries for others to reuse and re-purpose.
@@
These are only some challenges that the generative UI community is grappling with that would need to be resolved in this evolving landscape of Generative AI technology. It would be valuable to prioritize responsible AI and copyright ethics to improve the design of Misty. Insights from this design approach, and its influence on user interaction with the tool would contribute to HCI and Responsible AI.
@@
Our implementation of Misty is based on simple UI screens with limited interactivity. During our user study, participants expressed interest in expanding Misty to conceptual blending of interactive prototypes. Interactive UI prototypes contain many dimensions that are relevant for conceptual blending: user interactions, animations, and screen states. Tools like CoCapture (Chen et al., 2021) have made progress towards blending interactivity, and Keyframer (Tseng et al., 2024) has investigated the use of Generative AI for authoring animations. However, blending interactivity continues to pose design challenges. Blending static screenshots and sketches are relatively simple, yet it is not obvious how to specify sections and aspects—for example, of a video—for blending. Design principles from research in programming by demonstration can provide a useful starting point (Cypher and Halbert, 1993).
@@
8. Limitations
8.1. User Study Limitations
Our study participants were frontend developers recruited from a single US-based technology company. Other roles, such as UI/UX designers, may have different experiences with conceptual blending using Misty than developers. US-centric participants may also limit our understanding of how different culture design preferences might affect usage patterns and UI outputs. The first-use study format and controlled task environment may not always reflect how Misty would be used in real-world scenarios. A longitudinal study of Misty being used in the wild can be a valuable future work to answer these questions.
8.2. System and Technical Limitations
Misty works with static, single-screen mobile UI. It does not support the conceptual blending of interactive prototypes and UI animations. Understanding multi-screen, interactive UI prototypes with AI has been an active research area (Wu et al., 2023a; Lu et al., 2024a; Zhang et al., 2025). UI motion design protocols such as Lottie Files (LottieFiles, 2024) also make conceptual blending of UI animations easier. We expect to build Misty on top of these work to extend its blending capabilities.
Misty only supports accessibility incidentally (Wobbrock et al., 2011; Apple Inc., 2024a; Stephanidis et al., 1998; Swearngin et al., 2024). Misty can incorporate accessibility design best practices (Apple Inc., 2024a)—for example, by including post-hoc checking mechanisms to ensure accessibility—such as the ones in DesignChecker (Huh and Pavel, 2024). Adopting an ability-based design approach, Misty can also generate blending outputs based on user abilities, to serve diverse users’ needs.
Misty also does not consider cross-cultural design (Plocher et al., 2021; Hofstede, 2009; Bourges-Waldegg and Scrivener, 1998; Kellogg and Thomas, 1993) in conceptual blending. Research has shown that users from different cultures have diverse preferences for AI capabilities, control, and connections (Ge et al., 2024). Cultural contexts also influence how AI models and users influence each other’s decision making (Lee and Joshi, 2020). Misty could adopt more culturally-aware AI models in generating blending outputs (Suzuki, 2023; Robinson, 2020). Co-designing and testing Misty with users from different cultures can also help us see how their preferences and desired outputs may differ (Heaton, 2002).
8.3. Responsible AI Considerations in UI Prototyping
Practitioners and researchers have been incorporating Responsible AI and AI ethics into technical systems (Amershi et al., 2019; Wright et al., 2020; Shneiderman, 2020; Wang et al., 2023a; Varanasi and Goyal, 2023; Ali et al., 2023). Systems and toolkits have been proposed in HCI research to support Responsible AI in design and development workflows (Munechika et al., 2022; Cabrera et al., 2019; Wexler et al., 2019; Wang et al., 2022).
The boundary of “ethical” and “unethical copying” in UI/UX is fluid. Much of today’s UI design and development builds upon established design patterns to prioritize usability over creativity (Talton et al., 2012; Lu et al., 2024a). Also, open-source communities promote re-using and re-purposing pre-built UI components (Bach and Twidale, 2010; Zilouchian Moghaddam et al., 2012; Apple Inc., 2024b; Shadcn, 2024).
Before AI adoption in design tools, we already witnessed increased homogenization of website design (Goree et al., 2021). AI-enabled prototyping workflows, like Misty, can amplify this trend (Anderson et al., 2024). Automated prototyping tools raise questions about authorship and intellectual property. They risk copyright infringement and can spread biased design choices. Guidelines and ethical oversight can ensure these tools support creative freedom and minimize harm.
In Misty, we evaluated the experience of conceptual blending to understand its impact on UI prototyping workflows. Misty does not currently have safeguards. Design opportunities exist for Misty to prioritize Responsible AI considerations. For example, researchers can visualize an AI model’s conceptual blending outputs, to help users understand the output distributions. Combined with visual analytics (Wexler et al., 2019; Munechika et al., 2022), this could help users discover and understand AI model biases. Furthermore, in-situ support similar to Farsight (Wang et al., 2024) could increase users’ Responsible AI awareness in prototyping and adopting blending outputs. Future research in this direction can lead to research insights for conceptual blending in UI prototyping.